### v2

@@ 8 Generalized network utilities
$\pi_{v}$ $=0\,,\quad\forall v\in{\mathcal{M}}:\operatorname{net}_{f}(v)>-x_{v}\mbox{ or }v\in V\setminus({\mathcal{M}}\cup t):\operatorname{net}_{f}(v)>0\,,$
@@
$\pi$ $\geq 0\,.$
In accordance with this lemma, we say that $f\in\mathbb{R}^{E}$ and $\pi\in\mathbb{R}^{V}$ form a fitting pair with respect to $x$, if $f$ is a feasible generalized flow with supply $x$, and $\pi$ satisfies (9). Note that this implies that $f$ and $\pi$ are primal and dual optimal solutions to $\operatorname{GenflowLP}(x)$. Further, $g\in\partial^{\wedge}u(x)$ if and only if $g$ is the restriction of $\pi$ to ${\mathcal{M}}$ for a fitting pair $(f,\pi)$ with respect to $x$.
For every arc $e=(v,w)\in E$, let us define the reverse arc $\overleftarrow{e}=(v,w)$ with gain factor $\gamma_{\overleftarrow{e}}=1/\gamma_{e}$. By sending $\alpha$ units of flow on $\overleftarrow{e}$ we mean decreasing the flow value on $e$ by $\gamma_{e}\alpha$. Given a fitting pair $(f,\pi)$ with respect to $x$, we define the auxiliary network $G^{\textrm{aux}}_{f,\pi}=(V,E^{\textrm{aux}}_{f,\pi})$ by
$E^{\textrm{aux}}_{f,\pi}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{e=(v,w)\in E\,:\,\gamma_{e}\pi_{w}-\pi_{v}=0\mbox{ and }f_{e}<h_{e}\right\}\cup\left\{\overleftarrow{e}\,:\,e\in E\,:\,\gamma_{e}\pi_{w}-\pi_{v}=0\mbox{ and }0<f_{e}\right\}\,.$
In the proof of Theorem 1.5, we use two flow subroutines in the generalized flow network. The first one starts with a fitting pair $(f,\pi)$ with respect to a demand vector $x$, and outputs a flow $f^{\prime}$ for a decreased demand vector $x^{\prime}$ such that the flow value $\operatorname{net}_{f^{\prime}}(t)$ is decreased to the smallest possible amount while maintaining $(f^{\prime},\pi)$ as a fitting pair.
Input: $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, and a fitting pair $(f,\pi)$ with respect to $x$
Output: A supply vector $0\leq x^{\prime}\leq x$ and a feasible generalized flow $f^{\prime}$ such that $(f^{\prime},\pi)$ is a fitting pair with respect to $x^{\prime}$, and subject to this, $\operatorname{net}_{f^{\prime}}(t)$ is minimal.
@@
Lemma 8.3.
Pullback-Flow can be formulated as a linear program. For the output flow $f^{\prime}$, the auxiliary graph $E^{\textrm{aux}}_{f^{\prime},\pi}$ does not contain any directed paths from the sink $t$ to $\operatorname{supp}(x^{\prime})$.
@@
When fixing $\pi$, the requirement that $(f^{\prime},\pi)$ be a fitting pair can be formulated by linear constraints on $f^{\prime}$. For example, for every arc $e=(v,w)$, if $\gamma_{e}\pi_{w}-\pi_{v}<0$ then we require $f^{\prime}_{e}=0$, and if $\gamma_{e}\pi_{w}-\pi_{v}>0$ then $f^{\prime}_{e}=h_{e}$; if $\gamma_{e}\pi_{w}-\pi_{v}=0$ then any value $0\leq f^{\prime}_{e}\leq h_{e}$ is allowed. For the second part, assume for a contradiction that $E^{\textrm{aux}}_{f^{\prime},\pi}$ contains a directed path $P$ from $t$ to a node $j\in{\mathcal{M}}$ with $x^{\prime}_{j}>0$. Then, we can send a small amount of flow along this path by maintaining the flow balance on every intermediate node, and decreasing $\operatorname{net}_{f^{\prime}}(t)$ and $x^{\prime}_{j}$ by positive amounts. This contradicts the minimality of $\operatorname{net}_{f^{\prime}}(t)$. ∎
Remark 8.4.
Pullback-Flow can be in fact formulated as a (classical) maximum flow problem. To see this, one can use the standard generalized flow relabelling technique, first introduced by Glover Glover and Klingman (1973), see also e.g., Goldberg et al. (1991), Olver and Végh (2020)). That is, we multiply the flow at node $v$ by $\pi_{v}$ and replacing $\gamma_{e}$ by $\gamma_{e}\pi_{w}/\pi_{v}$ for every $e=(v,w)\in E$. On every arc, we fix $f^{\prime}_{e}$ to the upper or lower bound unless $\gamma_{e}\pi_{w}-\pi_{v}=0$, in which case the relabelled gain factor is 1.
Our second subroutine aims to increase the supply at a designated node $k\in{\mathcal{M}}$ while the supply at all other goods is non-increasing, and the flow value $\operatorname{net}_{f}(t)$ is non-decreasing. The goal is again to maintain a fitting pair $(f^{\prime},\pi)$ for a demand vector $x^{\prime}$.
Input: $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, and a fitting pair $(f,\pi)$ with respect to $x$, and a good $k\in{\mathcal{M}}$ such that $\pi_{k}>0$.
Output: A supply vector $x^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $x^{\prime}_{k}\geq x_{k}$ and $x^{\prime}_{j}\leq x_{j}$ if $j\in{\mathcal{M}}\setminus\{k\}$, and a feasible generalized flow $f^{\prime}$ such that $(f^{\prime},\pi)$ is a fitting pair with respect to $x^{\prime}$, $\operatorname{net}_{f}(t^{\prime})\geq\operatorname{net}_{f}(t)$, and subject to this, $x^{\prime}_{k}$ is maximal.
@@
The following lemma can be derived similarly to Lemma 8.3.
Lemma 8.5.
Reroute-Flow can be formulated as a linear program. For the output flow $f^{\prime}$, the auxiliary graph $E^{\textrm{aux}}_{f^{\prime},\pi}$ does not contain any directed path from $k$ to the set $(\operatorname{supp}(x^{\prime})\setminus\{k\})\cup\{t\}$.
@@
Lemma 8.6.
Consider a generalized network utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and $x\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$. Then, $x\in\mathcal{GD}^{u}(b,\mathbf{q})$ if and only if there exists a fitting pair $(f,\pi)$ with respect to $x$ such that $u(x)=\operatorname{net}_{f}(t)$ and
$\pi_{j}\leq\frac{q_{j}}{b}\cdot u(x)\,\quad\forall j\in{\mathcal{M}}\,,\,\text{ with equality for }j\in\operatorname{supp}(x)\,.$ (10)
Theorem 1.5 follows by showing the two properties in Definition 2.14 in Lemmas 8.7 and Lemma 8.8 below. We start by property (ii).
Lemma 8.7.
Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function, $b>b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and let $x\in\mathcal{GD}^{u}(b,\mathbf{q})$. Then, there exists $x^{\prime}\in\mathcal{GD}^{u}(b^{\prime},\mathbf{q})$ such that $x^{\prime}\leq x$.
@@
The proof is given by a finite algorithm, shown as Algorithm 3. We start from fitting pair $(f,\pi)$ with respect to $x$ as in Lemma 8.6. Thus, the inequality (10) holds for $\pi$, $x$, $\mathbf{q}$ and $b$. If $u(x)=0$, then we must have $\pi_{j}=0$ for all $j\in{\mathcal{M}}$ and thus (10) holds also when replacing $b$ by $b^{\prime}$, and therefore we can return $x^{\prime}=x$. For the rest of the proof, let us assume $u(x)>0$. Similarly, we can assume that $\mathbf{q}\neq\mathbf{0}$. Note that $\pi_{j}>0$ for every $j\in\operatorname{supp}(x)\cap\operatorname{supp}(\mathbf{q})$. When replacing $b$ by $b^{\prime}$, all inequalities with in (10) with $q_{j}>0$ become strict; the ratio of the right and left hand sides will be $\alpha\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}b/b^{\prime}>1$ for all $j\in\operatorname{supp}(x)\cap\operatorname{supp}(\mathbf{q})$. Also, recall $u(x)=\operatorname{net}_{f}(t)$.
Algorithm 3 alternates between two operations: changing $x$ and $f$ while keeping $\pi$ unchanged; and changing the dual $\pi$ while keeping $x$ and $f$ unchanged. Throughout, it maintains a fitting pair with respect to a decreasing vector $x^{\prime}$, such that the required equalities in (10) are gradually restored.
[…]
@@
If $\alpha^{\prime}\leq 1$, then we choose $\lambda\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}(\alpha-1)/(\alpha-\alpha^{\prime})\in[0,1]$ such that $\lambda\alpha+(1-\lambda)\alpha^{\prime}=1$. Then, $x^{\prime\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\lambda x^{\prime}+(1-\lambda)x$ and the flow $f^{\prime\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\lambda f^{\prime}+(1-\lambda)f$ satisfy that $(f^{\prime\prime},\pi)$ is a fitting pair with respect to $x^{\prime\prime}$, and (10) holds for $\pi$, $x^{\prime\prime}$ and $b^{\prime}$. Thus, we can terminate by outputting $x^{\prime\prime}\in\mathcal{GD}^{u}(b^{\prime},\mathbf{q})$.
@@
If $\alpha^{\prime}>1$, then we update the dual $\pi$ to $\pi^{\prime}$ as follows. Let $T\supseteq\operatorname{supp}(x^{\prime})$ be the set of nodes that can reach $\operatorname{supp}(x^{\prime})$ in the auxiliary graph $(G,E^{\textrm{aux}}_{f^{\prime},\pi})$. By Lemma 8.3, $t\notin T$. Let us define $\pi^{\prime}$ such that $\pi^{\prime}_{v}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\pi_{v}$ for $v\notin T$ and $\pi^{\prime}_{v}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\delta\pi_{v}$ for $v\in T$ for the largest value $1\leq\delta\leq\alpha^{\prime}$ such that $(f^{\prime},\pi^{\prime})$ is a fitting pair.

### v3

@@ Appendix D Generalized network utilities
$A\mathbf{z}$ $\leq B\mathbf{x}+\mathbf{b}$
$\mathbf{z}$ $\geq\mathbf{0}$
Then, for every $\mathbf{x}\in\mathcal{D}$, $\mathbf{g}\in\partial^{\wedge}\varphi(\mathbf{x})$ if and only if $\mathbf{g}=B^{\top}\bm{\pi}$ for an optimal solution $\bm{\pi}\in\mathbb{R}_{\geq 0}^{k}$ to the dual linear program
[…]
$A^{\top}\bm{\pi}$ $\geq\mathbf{c}$
$\bm{\pi}$ $\geq\mathbf{0}\,.$
Proof.
First, assume that $\bm{\pi}\in\mathbb{R}_{\geq 0}^{\ell}$ is an optimal solution to $\operatorname{Dual}(\mathbf{x})$ for a given $\mathbf{x}\in\mathbb{R}^{m}$. We need to show that $\mathbf{g}=B^{\top}\bm{\pi}$ is a supergradient at $\mathbf{x}$, that is, $\varphi(\mathbf{t})\leq\varphi(\mathbf{x})+\left\langle\mathbf{g},\mathbf{t}-\mathbf{x}\right\rangle$ for any $\mathbf{t}\in\mathcal{D}$.
The programs $\operatorname{Dual}(\mathbf{x})$ and $\operatorname{Dual}(\mathbf{t})$ have the same feasible region, with linear objective functions $B\mathbf{x}+\mathbf{b}$ and $B\mathbf{t}+\mathbf{b}$, respectively. By weak duality, the cost of every feasible dual solution is at least the primal optimum value, thus, $\varphi(\mathbf{t})\leq\mathbf{t}^{\top}B\bm{\pi}+\mathbf{b}^{\top}\bm{\pi}=\mathbf{g}^{\top}\mathbf{t}+\mathbf{b}^{\top}\bm{\pi}$. The statement follows as
$\varphi(\mathbf{t})\leq\mathbf{g}^{\top}\mathbf{t}+\mathbf{b}^{\top}\bm{\pi}=\mathbf{g}^{\top}(\mathbf{t}-\mathbf{x})+\mathbf{x}^{\top}B\bm{\pi}+\mathbf{b}^{\top}\bm{\pi}=\varphi(\mathbf{x})+\left\langle\mathbf{g},\mathbf{t}-\mathbf{x}\right\rangle\,$,
noting also that $\varphi(\mathbf{x})=\mathbf{x}^{\top}B\bm{\pi}+\mathbf{b}^{\top}\bm{\pi}$ since $\bm{\pi}$ is an optimal solution to $\operatorname{Dual}(\mathbf{x})$, and $\varphi(\mathbf{x})$ is the optimum value.
Conversely, we need to show that if $\mathbf{g}\in\partial^{\wedge}\varphi(\mathbf{x})$, then there exists a dual optimal solution $\bm{\pi}$ such that $\mathbf{g}=B^{\top}\bm{\pi}$. Let $\mathbf{z}$ be a primal optimal solution to $\operatorname{Primal}(\mathbf{x})$. Using complementary slackness, we can turn $\operatorname{Dual}(\mathbf{x})$ into a feasibility system. Namely, $\mathbf{g}=B^{\top}\bm{\pi}$ for a dual optimal solution if and only if
$B^{\top}\bm{\pi}$ $=\mathbf{g}$ (9)
$A_{i}^{\top}\bm{\pi}$ $=c_{i}\,,\quad\forall i:z_{i}>0$
$A_{i}^{\top}\bm{\pi}$ $\geq c_{i}\,,\quad\forall i:z_{i}=0$
$\pi_{j}$ $=0\,,\quad\forall j:(A\mathbf{z})_{j}<(B\mathbf{x}+\mathbf{b})_{j}$
$\pi_{j}$ $\geq 0,\quad\forall j:(A\mathbf{z})_{j}=(B\mathbf{x}+\mathbf{b})_{j}\,.$
Here, $A_{i}$ denotes the $i$-th column of $A$. By Farkas’s lemma, if (9) is not feasible, then the following system has a feasible solution:
$(A\mathbf{w})_{j}+(B\mathbf{r})_{j}$ $\leq 0\,,\quad\forall j:(A\mathbf{z})_{j}=(B\mathbf{x}+\mathbf{b})_{j}$
$w_{i}$ $\geq 0\,,\quad\forall i:z_{i}=0$
$\mathbf{c}^{\top}\mathbf{w}+\mathbf{g}^{\top}\mathbf{r}$ $>0\,.$
For a small enough $\varepsilon>0$, $\mathbf{z}+\varepsilon\mathbf{w}$ is feasible to the system $\operatorname{Primal}(\mathbf{x}-\varepsilon\mathbf{r})$. Hence, for $\mathbf{t}=\mathbf{x}-\varepsilon\mathbf{r}$, $\varphi(\mathbf{t})\geq\mathbf{c}^{\top}(\mathbf{z}+\varepsilon\mathbf{w})=\mathbf{c}^{\top}\mathbf{z}+\varepsilon\mathbf{c}^{\top}\mathbf{w}$. Using $\mathbf{c}^{\top}\mathbf{z}=\varphi(\mathbf{x})$ by the choice of $\mathbf{z}$ as an optimal solution to $\operatorname{Primal}(\mathbf{x})$, and the bound $\mathbf{c}^{\top}\mathbf{w}>-\mathbf{g}^{\top}\mathbf{r}$, we get $\varphi(\mathbf{t})>\varphi(\mathbf{x})-\varepsilon\mathbf{g}^{\top}\mathbf{r}=\varphi(\mathbf{x})+\left\langle\mathbf{g},\mathbf{t}-\mathbf{x}\right\rangle$, a contradiction to $\mathbf{g}\in\partial^{\wedge}\varphi(\mathbf{x})$. ∎
Consider a generalized flow instance defined in a graph $G=(V,E)$, sink $t\in V$, capacities $\mathbf{h}\in\mathbb{R}_{\geq 0}^{E}$, gain factors $\bm{\gamma}\in\mathbb{R}_{>0}^{E}$ and supply set $S={\mathcal{M}}$. For the rest of this section, let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function defined in the form $\operatorname{GenflowLP}(\mathbf{x})$ in Section 2.
Using Lemma D.1 and complementary slackness, we get the following characterization of subgradients of generalized network utility functions. Note that for a generalized network utility function, the domain is $\mathcal{D}=\mathbb{R}_{\geq 0}^{\mathcal{M}}$, because $\mathbf{f}=0$ is always a feasible solution to $\operatorname{GenflowLP}(\mathbf{x})$, and the definition assumes that the objective is always bounded.
Lemma D.2.
Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{+}$ be a generalized network utility function defined in the form $\operatorname{GenflowLP}(\mathbf{x})$ for $S={\mathcal{M}}$. For a given supply vector $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, let $\mathbf{f}$ be an optimal solution to $\operatorname{GenflowLP}(\mathbf{x})$. Then, $\mathbf{g}\in\partial^{\wedge}u(\mathbf{x})$ if and only if $\mathbf{g}=\bm{\pi}|_{\mathcal{M}}$ for a vector $\bm{\pi}\in\mathbb{R}_{\geq 0}^{V}$ that satisfies
@@
$\pi_{v}$ $=0\,,\quad\forall v\in{\mathcal{M}}:\operatorname{net}_{\mathbf{f}}(v)>-x_{v}\mbox{ or }v\in V\setminus({\mathcal{M}}\cup t):\operatorname{net}_{\mathbf{f}}(v)>0\,$,
@@
$\bm{\pi}$ $\geq\mathbf{0}\,.$
In accordance with this lemma, we say that $\mathbf{f}\in\mathbb{R}^{E}$ and $\bm{\pi}\in\mathbb{R}^{V}$ form a fitting pair with respect to $\mathbf{x}$, if $\mathbf{f}$ is a feasible generalized flow with supply $\mathbf{x}$, and $\bm{\pi}$ satisfies (10). Note that this implies that $\mathbf{f}$ and $\bm{\pi}$ are primal and dual optimal solutions to $\operatorname{GenflowLP}(\mathbf{x})$. Further, $\mathbf{g}\in\partial^{\wedge}u(\mathbf{x})$ if and only if $\mathbf{g}$ is the restriction of $\bm{\pi}$ to $\mathcal{M}$} for a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$.
For every arc $e=(v,w)\in E$, let us define the reverse arc $\overleftarrow{e}=(v,w)$ with gain factor $\gamma_{\overleftarrow{e}}=1/\gamma_{e}$. By sending $\alpha$ units of flow on $\overleftarrow{e}$ we mean decreasing the flow value on $e$ by $\gamma_{e}\alpha$. Given a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$, we define the auxiliary network $G^{\textrm{aux}}_{\mathbf{f},\bm{\pi}}=(V,E^{\textrm{aux}}_{\mathbf{f},\bm{\pi}})$ by
$E^{\textrm{aux}}_{\mathbf{f},\bm{\pi}}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\left\{e=(v,w)\in E\,:\,\gamma_{e}\pi_{w}-\pi_{v}=0\mbox{ and }f_{e}<h_{e}\right\}\cup\left\{\overleftarrow{e}\,:\,e\in E\,:\,\gamma_{e}\pi_{w}-\pi_{v}=0\mbox{ and }0<f_{e}\right\}\,.$
In the proof of Theorem 1.5, we use two flow subroutines in the generalized flow network. The first one starts with a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to a demand vector $\mathbf{x}$, and outputs a flow $\mathbf{f}^{\prime}$ for a decreased demand vector $\mathbf{x}^{\prime}$ such that the flow value $\operatorname{net}_{\mathbf{f}^{\prime}}(t)$ is decreased to the smallest possible amount while maintaining $(\mathbf{f}^{\prime},\bm{\pi})$ as a fitting pair.
Input: $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, and a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$
Output: A supply vector $\mathbf{0}\leq\mathbf{x}^{\prime}\leq\mathbf{x}$ and a feasible generalized flow $\mathbf{f}^{\prime}$ such that $(\mathbf{f}^{\prime},\bm{\pi})$ is a fitting pair with respect to $\mathbf{x}^{\prime}$, and subject to this, $\operatorname{net}_{\mathbf{f}^{\prime}}(t)$ is minimal.
@@
Lemma D.3.
Pullback-Flow can be formulated as a linear program. For the output flow $\mathbf{f}^{\prime}$, the auxiliary graph $E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$ does not contain any directed paths from the sink $t$ to $\operatorname{supp}(\mathbf{x}^{\prime})$.
@@
When fixing $\bm{\pi}$, the requirement that $(\mathbf{f}^{\prime},\bm{\pi})$ be a fitting pair can be formulated by linear constraints on $\mathbf{f}^{\prime}$. For example, for every arc $e=(v,w)$, if $\gamma_{e}\pi_{w}-\pi_{v}<0$ then we require $f^{\prime}_{e}=0$, and if $\gamma_{e}\pi_{w}-\pi_{v}>0$ then $f^{\prime}_{e}=h_{e}$; if $\gamma_{e}\pi_{w}-\pi_{v}=0$ then any value $0\leq f^{\prime}_{e}\leq h_{e}$ is allowed. For the second part, assume for a contradiction that $E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$ contains a directed path $P$ from $t$ to a node $j\in{\mathcal{M}}$ with $x^{\prime}_{j}>0$. Then, we can send a small amount of flow along this path by maintaining the flow balance on every intermediate node, and decreasing $\operatorname{net}_{\mathbf{f}^{\prime}}(t)$ and $x^{\prime}_{j}$ by positive amounts. This contradicts the minimality of $\operatorname{net}_{\mathbf{f}^{\prime}}(t)$. ∎
Remark D.4.
Pullback-Flow can be in fact formulated as a (classical) maximum flow problem. To see this, one can use the standard generalized flow relabeling technique, first introduced by Glover Glover and Klingman (1973), see also e.g., Goldberg et al. (1991), Olver and Végh (2020)). That is, we multiply the flow at node $v$ by $\pi_{v}$ and replacing $\gamma_{e}$ by $\gamma_{e}\pi_{w}/\pi_{v}$ for every $e=(v,w)\in E$. On every arc, we fix $f^{\prime}_{e}$ to the upper or lower bound unless $\gamma_{e}\pi_{w}-\pi_{v}=0$, in which case the relabeled gain factor is 1.
Our second subroutine aims to increase the supply at a designated node $k\in{\mathcal{M}}$ while the supply at all other goods is non-increasing, and the flow value $\operatorname{net}_{\mathbf{f}}(t)$ is non-decreasing. The goal is again to maintain a fitting pair $(\mathbf{f}^{\prime},\bm{\pi})$ for a demand vector $\mathbf{x}^{\prime}$.
Input: $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, and a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$, and a good $k\in{\mathcal{M}}$ such that $\pi_{k}>0$.
Output: A supply vector $\mathbf{x}^{\prime}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $x^{\prime}_{k}\geq x_{k}$ and $x^{\prime}_{j}\leq x_{j}$ if $j\in{\mathcal{M}}\setminus\{k\}$, and a feasible generalized flow $\mathbf{f}^{\prime}$ such that $(\mathbf{f}^{\prime},\bm{\pi})$ is a fitting pair with respect to $\mathbf{x}^{\prime}$, $\operatorname{net}_{\mathbf{f}}(t^{\prime})\geq\operatorname{net}_{\mathbf{f}}(t)$, and subject to this, $x^{\prime}_{k}$ is maximal.
@@
The following lemma can be derived similarly to Lemma D.3.
Lemma D.5.
Reroute-Flow can be formulated as a linear program. For the output flow $\mathbf{f}^{\prime}$, the auxiliary graph $E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$ does not contain any directed path from $k$ to the set $(\operatorname{supp}(\mathbf{x}^{\prime})\setminus\{k\})\cup\{t\}$.
@@
Lemma D.6.
Consider a generalized network utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and $\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$. Then, $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{q},b)$ if and only if there exists a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$ such that $u(\mathbf{x})=\operatorname{net}_{\mathbf{f}}(t)$ and
$\pi_{j}\leq\frac{q_{j}}{b}\cdot u(\mathbf{x})\,\quad\forall j\in{\mathcal{M}}\,,\,\text{ with equality for }j\in\operatorname{supp}(\mathbf{x})\,.$ (11)
Theorem 1.5 follows by showing the two properties in Definition 2.14 in Lemmas D.7 and Lemma D.8 below. We start by property (ii).
Lemma D.7.
Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function, $b>b^{\prime}>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, and let $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{q},b)$. Then, there exists $\mathbf{x}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$ such that $\mathbf{x}^{\prime}\leq\mathbf{x}$.
@@
The proof is given by a finite algorithm, shown as Algorithm 3. We start from fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$ as in Lemma D.6. Thus, the inequality (11) holds for $\bm{\pi}$, $\mathbf{x}$, $\mathbf{q}$ and $b$. If $u(\mathbf{x})=0$, then we must have $\pi_{j}=0$ for all $j\in{\mathcal{M}}$ and thus (11) holds also when replacing $b$ by $b^{\prime}$, and therefore we can return $\mathbf{x}^{\prime}=\mathbf{x}$. For the rest of the proof, let us assume $u(\mathbf{x})>0$. Similarly, we can assume that $\mathbf{q}\neq\mathbf{0}$. Note that $\pi_{j}>0$ for every $j\in\operatorname{supp}(\mathbf{x})\cap\operatorname{supp}(\mathbf{q})$. When replacing $b$ by $b^{\prime}$, all inequalities with in (11) with $q_{j}>0$ become strict; the ratio of the right and left hand sides will be $\alpha\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}b/b^{\prime}>1$ for all $j\in\operatorname{supp}(\mathbf{x})\cap\operatorname{supp}(\mathbf{q})$. Also, recall $u(\mathbf{x})=\operatorname{net}_{\mathbf{f}}(t)$.
Algorithm 3 alternates between two operations: changing $\mathbf{x}$ and $\mathbf{f}$ while keeping $\bm{\pi}$ unchanged; and changing the dual $\bm{\pi}$ while keeping $\mathbf{x}$ and $\mathbf{f}$ unchanged. Throughout, it maintains a fitting pair with respect to a decreasing vector $\mathbf{x}^{\prime}$, such that the required equalities in (11) are gradually restored.
In the first step, we call $\textnormal{{Pullback-Flow}}(\mathbf{x},\mathbf{f},\bm{\pi})$ to get demands $\mathbf{x}^{\prime}\leq\mathbf{x}$ and a corresponding fitting pair $(\mathbf{f}^{\prime},\bm{\pi})$. We have $u(\mathbf{x}^{\prime})=\operatorname{net}_{\mathbf{f}^{\prime}}(t)\leq\operatorname{net}_{\mathbf{f}}(t)=u(\mathbf{x})$. Note that $\frac{q_{j}}{b^{\prime}}u(\mathbf{x}^{\prime})/\pi_{j}$ is the same value $\alpha^{\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\alpha u(\mathbf{x}^{\prime})/u(\mathbf{x})$ for every $j\in\operatorname{supp}(\mathbf{x}^{\prime})\cap\operatorname{supp}(\mathbf{q})$.
@@
If $\alpha^{\prime}\leq 1$, then we choose $\lambda\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}(\alpha-1)/(\alpha-\alpha^{\prime})\in[0,1]$ such that $\lambda\alpha+(1-\lambda)\alpha^{\prime}=1$. Then, $\mathbf{x}^{\prime\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\lambda\mathbf{x}^{\prime}+(1-\lambda)\mathbf{x}$ and the flow $\mathbf{f}^{\prime\prime}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\lambda\mathbf{f}^{\prime}+(1-\lambda)\mathbf{f}$ satisfy that $(\mathbf{f}^{\prime\prime},\bm{\pi})$ is a fitting pair with respect to $\mathbf{x}^{\prime\prime}$, and (11) holds for $\bm{\pi}$, $\mathbf{x}^{\prime\prime}$ and $b^{\prime}$. Thus, we can terminate by outputting $\mathbf{x}^{\prime\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$.
@@
If $\alpha^{\prime}>1$, then we update the dual $\bm{\pi}$ to $\bm{\pi}^{\prime}$ as follows. Let $T\supseteq\operatorname{supp}(\mathbf{x}^{\prime})$ be the set of nodes that can reach $\operatorname{supp}(\mathbf{x}^{\prime})$ in the auxiliary graph $(G,E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}})$. By Lemma D.3, $t\notin T$. Let us define $\bm{\pi}^{\prime}$ such that $\pi^{\prime}_{v}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\pi_{v}$ for $v\notin T$ and $\pi^{\prime}_{v}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\delta^{\prime}\pi_{v}$ for $v\in T$ for the largest value $1\leq\delta^{\prime}\leq\alpha^{\prime}$ such that $(\mathbf{f}^{\prime},\bm{\pi}^{\prime})$ is a fitting pair.